OpenStudy (samara8954):

A scientist estimated that a mixture would need 3 milliliters of a chemical to balance. The actual amount needed was 7 milliliters. What was the percent error of the scientist's estimation? 133.33% 3% 42.86% 57.14%

OpenStudy (samara8954):

@mathstudent55

OpenStudy (mokeira):

find the error (7-3)= error then change to precentage

OpenStudy (mathstudent55):

\(percent ~error = \dfrac{new~value - old~value}{old~value} \times 100\%\) Use old value = 3 and new value = .

OpenStudy (samara8954):

so the precentage is 3

OpenStudy (samara8954):

am I correct

OpenStudy (samara8954):

so now what

OpenStudy (mathstudent55):

Wait. I just reread the question. I made a mistake. The formula above is correct. You need to use 7 as the old value since 7 is what we are comparing to. 3 is the new value since this is the value whose percent error we want. \(percent ~error = \dfrac{new~value - old~value}{old~value} \times 100\%\) \(percent ~error = \dfrac{3 - 7}{7} \times 100\%\)

OpenStudy (mathstudent55):

\(\dfrac{3 - 7}{7} \times 100\% = \dfrac{-4}{7} \times 100\% = -0.5714 \times 100\% = -57.14\%\) The error in the estimate is 57.14%. The negative sign in the formula means the the estimate is 57.14% lower than the correct value.